OpenStudy (anonymous):

Just another cute problem, How many eight digit numbers formed using the digit \( \{7,8,9\} \) has exactly three \( 7 \)'s ?

OpenStudy (anonymous):

5P2*3P1?

OpenStudy (anonymous):

Well, Ishaan it's one of my cute problem so may be not that easy ;-)

OpenStudy (anonymous):

Lol, Okay. I will try again.

OpenStudy (anonymous):

sure :D

OpenStudy (anonymous):

Btw Ishaan, I think this problem is IIT JEE standard, and might be extended to any undergrad level entrance test, so good luck! :) You can do it.

OpenStudy (anonymous):

\[--------\] 8 slots are to be filled by three numbers {7,8,9}, we need numbers who have exactly 3 7s. Implies, 5 slots left for {8,9}, order matters, 5P2 must be the number of ways slots can be filled. 3 slots left, and I am supposed to fill them up with 7s, order doesn't matter, so only 1 way. 5P2, is the answer 20?

OpenStudy (anonymous):

I did something wrong, I need to get hold on Combinatorics.

OpenStudy (anonymous):

Oh wait, the choice of filling 5slots with numbers 8 and 9, depends upon the arrangement of 3 slots (which are to be filled by 7s). So, 8C3*5P2. I am not sure again, just trying to improvise.

OpenStudy (zarkon):

\[2^{5}\cdot{8\choose 3}\]

OpenStudy (anonymous):

That's gives the correct answer @Zarkon Actually my approach is a bit mildly tedious so you may like to explain your approach for us mortal.

OpenStudy (zarkon):

if there are 8 digits and 3 of them must be 7's then 5 must be 8's or 9's so for each of those 5 numbers there are 2 choices. thus \(2^5\) options. now we just choose the 3 places to put the sevens. out of the 8 possible position we choose 3.

OpenStudy (anonymous):

But 2^{5} would include repetition of the numbers, wouldn't it?

OpenStudy (zarkon):

no...there are only 32 (not including the 7's) of them...list them out to convince yourself.

OpenStudy (anonymous):

Oh, thanks zarkon.

OpenStudy (anonymous):

My solution is : \[ \Huge \sum \limits_{n= 0}^5 \frac{8!}{3! \times n! \times {(5-n)!}} \]

OpenStudy (zarkon):

\[ \sum \limits_{n= 0}^5 \frac{8!}{3! \times n! \times {(5-n)!}}\] \[=\sum \limits_{n= 0}^5 \frac{8!\times 5!}{3!\times 5! \times n! \times {(5-n)!}}\] \[=\frac{8!}{3!\times 5!}\sum \limits_{n= 0}^5 \frac{5!}{n! \times {(5-n)!}}\] \[={8\choose 3}\sum \limits_{n= 0}^5 \frac{5!}{n! \times {(5-n)!}}\] \[={8\choose 3}\sum \limits_{n= 0}^5 \frac{5!}{n! \times {(5-n)!}}1^n1^{5-n}\] \[={8\choose 3}(1+1)^5\] \[={8\choose 3}\cdot2^5\]

OpenStudy (anonymous):

@Zarkon: Great answer, I like the way you reduced \( \large \sum \limits_{n= 0}^5 \frac{5!}{n! \times {(5-n)!}} \; to \; 2^5 \) using binomial theorem!

OpenStudy (zarkon):

yep..cute problem.

OpenStudy (anonymous):

Something happened to latex, is it rendering on server side now?

OpenStudy (anonymous):

Cutie pie :D btw we could also do this \[ \large \sum \limits_{n= 0}^5 \frac{5!}{n! \times {(5-n)!}} = \large \sum \limits_{n= 0}^5\binom{5}{n} = 2^5\] Using the identity, \(\large \sum \limits_{r= 0}^n\binom{n}{r} = 2^n \)

OpenStudy (anonymous):

Just to triger the notification system @Zarkon :-)

OpenStudy (anonymous):

@Ishaan94: No problem in my side.

OpenStudy (zarkon):

which again...is really the binomial theorem

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

We just need to expand \( (1+x)^n\) using binomial theorem and substitute \(x=1\) that will gives us the desired identity.